in reaction below, calculate delta G0 and Keq at 25 C (use values from the table for standrad free energies or hydrolysis)

ATP + pyruvate --> phosphoenolpyruvate + ADP
Answer #1

Hydrolysis Of ATP

ATPADP

∆G°=-30.5kJ/ mol

Hydrolysis of phosphoenolpyruvate

Phosphoenolpyruvatepyurvate

∆G°=-62kJ/mol

The reaction is:

ATP + PYURVATE ADP + PHOSPHOENOLPYRUVATE

∆G°= -30.5+62=31.5KJ/mol = 31.5J/ mol

Also,

∆G°= - RT ln K

∆G°= - 2.303 RT log K

Log K = -31.5×10^3/8.314×298×2.303=-5.520

K=antilog(-5.520?

K= 3.01×10^-6

Where K is equilibrium constant.